Find equations of the tangent lines to the curve

y = (x-1)/(x+1)

that are parallel to the line x − 2y = 5.
Answer #1

The tangent line is parallel to the line x - 2y = 5, this means their slopes will be same.

So, slope of the tangent line = 1/2

Given : y= r+1 dy- x +1-(x-1 x+1 )- r 1 2 or +2x-3-0 so, (r-1)(x 3)0 For x 1, y 0 or x=-5, y So, equation of tan gents: x-2y-